# TopoX: A Suite of Python Packages for Machine Learning on Topological Domains

### 04 Feb 2024 v1

@@ Abstract
We introduce TopoX, a Python software suite that provides reliable and user-friendly building blocks for computing and machine learning on topological domains that extend graphs: hypergraphs, simplicial, cellular, path and combinatorial complexes. TopoX consists of three packages: TopoNetX facilitates constructing and computing on these domains, including working with nodes, edges and higher-order cells; TopoEmbedX provides methods to embed topological domains into vector spaces, akin to popular graph-based embedding algorithms such as node2vec; TopoModelX is built on top of PyTorch and offers a comprehensive toolbox of higher-order message passing functions for neural networks on topological domains. The extensively documented and unit-tested source code of TopoX is available under MIT license at https://github.com/pyt-team.
@@ 1 Introduction
Deep learning has traditionally focused on processing data within Euclidean domains, as exemplified by the development of convolutional neural networks (CNNs) (Krizhevsky et al., 2012) and recurrent neural networks (RNNs) (Sutskever et al., 2014; Bahdanau et al., 2014). These models have been tailored to handle data supported on regular, grid-like structures present in images or sequences, capitalizing on information inherent within Euclidean domains. To process data supported on more diverse domains, geometric deep learning (GDL) has recently emerged as a powerful paradigm, extending the capabilities of deep learning to non-Euclidean data (Bronstein et al., 2021; Zhou et al., 2020; Cao et al., 2020). GDL leverages principles of geometric regularity, such as symmetries, invariances, or equivariances, introducing suitable inductive biases when processing diverse data domains, including sets, grids, manifolds, and graphs. In this context, Topological Deep Learning (TDL) (Hajij et al., ) represents a new, burgeoning frontier in the field of machine learning. TDL explores the development of models that go beyond traditional graph-based abstractions and provides a principled approach to process data that reside in domains characterized by multi-way relations such as simplicial complexes, cell complexes, and hypergraphs. These extensions allow for the representation of diverse data domains encountered in scientific computations. (Feng et al., 2019; Schaub et al., 2020; Hajij et al., 2020; Schaub et al., 2021; Roddenberry et al., 2021; Giusti et al., 2022; Yang and Isufi, 2023; Barbarossa and Sardellitti, 2020). While there have been significant advancements in the theory of topological deep learning, its practical implementation has faced challenges due to the absence of accessible software libraries that simplify computations and support deep learning models using higher-order structures.
@@ 3 Comparison and Interaction with Other Packages
The closest package to TopoEmbedX is KarateClub (Rozemberczki et al., 2020), which is a Python package consisting of methods for unsupervised learning on graph-structured data. TopoEmbedX extends the functionality of KarateClub to topological domains supported in TopoNetX.
PyG and DGL (Deep Graph Library) (Wang et al., 2019), which are two popular geometric deep learning packages that support deep learning models on graphs, are closely related to TopoModelX. The DHG (Deep HyperGraph) package (Feng et al., 2019) supports deep learning models defined on hypergraphs.
All three of our packages feature a continuous integration pipeline and are well-tested with code coverage of >=95% per package. This is comparable or better than many related graph-based learning libraries, such as PyG, DGL, XGI, NetworkX, and HyperNetX.

### 17 Feb 2024 v4

@@ Abstract
We introduce TopoX, a Python software suite that provides reliable and user-friendly building blocks for computing and machine learning on topological domains that extend graphs: hypergraphs, simplicial, cellular, path and combinatorial complexes. TopoX consists of three packages: TopoNetX facilitates constructing and computing on these domains, including working with nodes, edges and higher-order cells; TopoEmbedX provides methods to embed topological domains into vector spaces, akin to popular graph-based embedding algorithms such as node2vec; TopoModelX is built on top of PyTorch and offers a comprehensive toolbox of higher-order message passing functions for neural networks on topological domains. The extensively documented and unit-tested source code of TopoX is available under MIT license at https://pyt-team.github.io/.
@@ 1 Introduction
Deep learning traditionally operates within Euclidean domains, focusing on structured data like images (Krizhevsky et al., 2012) and sequences (Sutskever et al., 2014). However, to handle more diverse data types, geometric deep learning (GDL) (Bronstein et al., 2021; Zhou et al., 2020; Cao et al., 2020) has emerged. GDL extends deep learning to non-Euclidean data by leveraging geometric regularities like symmetries and invariances. Recently, Topological Deep Learning (TDL) (Hajij et al., ) has gained attention, exploring models beyond traditional graph-based abstractions to process data with multi-way relations, such as simplicial complexes and hypergraphs. These extensions allow for the representation of diverse data domains encountered in scientific computations (Feng et al., 2019; Schaub et al., 2020; Hajij et al., 2020; Schaub et al., 2021; Roddenberry et al., 2021; Giusti et al., 2022; Yang and Isufi, 2023; Barbarossa and Sardellitti, 2020). Despite theoretical advancements, practical implementation faces challenges due to the lack of accessible software libraries supporting deep learning models with higher-order structures.
@@ 3 Comparison and Interaction with Other Packages
The closest package to TopoEmbedX is KarateClub (Rozemberczki et al., 2020), which is a Python package consisting of methods for unsupervised learning on graph-structured data. TopoEmbedX extends the functionality of KarateClub to topological domains supported in TopoNetX. PyG and DGL (Deep Graph Library) (Wang et al., 2019), which are two popular geometric deep learning packages that support deep learning models on graphs, are closely related to TopoModelX. The DHG (Deep HyperGraph) package (Feng et al., 2019) supports deep learning models defined on hypergraphs. All three of our packages feature a continuous integration pipeline and are well-tested with code coverage of >=95% per package. This is comparable or better than many related graph-based learning libraries, such as PyG, DGL, XGI, NetworkX, and HyperNetX.